Perform each of the following calculations and exp
Answer #1

a) =0.6274×1.00×103/36.13       (36.13347612.... rounds upto 36.13 as 2.205 has 4 significant figures)

    =0.02×103 =2.0×10                       (0.017365...... rounds upto 2 significant figures)

Note: Exact number has infinite number of significant figures

b) = 70×10-15 (72 rounds upto one significant figure )

c) = 5.70×10-23    ( 5.701268....... rounds upto 3 significant figures)

d) =35.8×10-4    (35.83735.......rounds upto 3 significant figures)

Note: Multiplication and Division- consider the lowest number of significant figures